OpenStudy (anonymous):

-2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3) Determine which system below will produce infinitely many solutions. -4x + 6y = 20 4x - 3y = -2 -2x + 4y = 18 4x - 8y = -36 3x - 2y = 5 -6x - 4y = -10 5x - 3y = 12 15x - 9y = 24

OpenStudy (anonymous):

Solve for y for one equation and substitute the y with what you got into the other equation.

OpenStudy (snowfire):

A system with infinitely many solutions is one that has all lines overlapping each other entirely. Find the system that has two equations that are actually just multiples of each other (This is how I would initially solve the problem, otherwise I'd resort to substitution as Grazes said).

OpenStudy (anonymous):

For the first question is it b?

OpenStudy (anonymous):

no its (1,3) srry i mistyped

OpenStudy (snowfire):

Oh right I forgot this was actually two questions ^^" Yes the first question is (1,3)

OpenStudy (anonymous):

Oh man i got it mixed up but thank you so much! can you help me with the second one too?

OpenStudy (anonymous):

hahaha its ok i got that one! (:

OpenStudy (snowfire):

My first post was my explanation to the second question.

OpenStudy (anonymous):

is it D?

OpenStudy (snowfire):

If D was contained two equations that are actually just multiples, then what number would you multiply the first by to get the second? Remember that the 12 and 24 are parts of this rule.

OpenStudy (snowfire):

If D contained*

OpenStudy (anonymous):

i think it"s B

OpenStudy (snowfire):

Well it is, but I wanted Veronica to work for it =P

OpenStudy (anonymous):

hahaha thank you guys so much! Can you help me with one more? Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - 2y = 6 2x + y = 5 -4x - 2y = 10 -4y = 4 -4x - 5y = -1 -7y = 5 3x + 2y = 6 7x = 12 4x + 2y = 10 8x = 16

OpenStudy (anonymous):

i was thinking it might be c? im confused on this one.

OpenStudy (anonymous):

is 4x - 2y = 6 the given solution ? 2x + y = 5

OpenStudy (snowfire):

Same here, I don't see where the solution is

OpenStudy (anonymous):

yes thats the given solution! 4x - 2y = 6 2x + y = 5

OpenStudy (anonymous):

the x and y for the original solution is (2,1) so find that one that gives u (2,1)

OpenStudy (snowfire):

You're basically doing the same thing as before, except this time you just find the solution to your given system, and apply that to each of the answer choices. Whichever system is true with it is the answer.

OpenStudy (anonymous):

alright im thinking it's A?

OpenStudy (snowfire):

Plugging in the solution (2,1) does not make A true.

OpenStudy (anonymous):

A is not true.. plug the points like snowfire said

OpenStudy (anonymous):

how do i plug in the points? can you show the correct way in to setting it up? then i will solve it?

OpenStudy (snowfire):

The 2 and 1 in (2,1) are the values for x and y respectively. Plug in 2 for every x and 1 for every y

OpenStudy (anonymous):

plug 2 for x and 1 for y 4*2 - 2*1 = 6 2*2 + 1 = 5

OpenStudy (anonymous):

C :)

OpenStudy (anonymous):

i'm not very good at math :/ haha thank you for being so patient with me

OpenStudy (anonymous):

it's okay, take ur time, and no it is not C

OpenStudy (anonymous):

alright, it's b

OpenStudy (anonymous):

dont just guess, try solving it

OpenStudy (anonymous):

does -7*1 gives u 5? it does not so it is not b

OpenStudy (anonymous):

OHHH!!! wait are you solving it by multiplying?

OpenStudy (anonymous):

yeah sub the point to x and y and solve them.. i gave u an example at the top

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

its d :)

OpenStudy (anonymous):

yup